if f(x) = cot ^-1 x^x -x^-x/22 then find dy/ dx  at x =1 

Grade:11

1 Answers

Niyati Parekh
5 years ago
The answer to the above question 
f(x)=cot^-1(x^x - x^-x/2) , f'(1)= -1
 
cot^-1(x)= -1/1+x^2
d(x^x)/dx = x^x(1+logx)